OpenStudy (lncognlto):

Solve the equation: sqr3 sinx = 1/2 sec x

OpenStudy (lncognlto):

\[\sqrt{3} \sin x = \frac{ 1 }{ 2 } \sec x\]

OpenStudy (lncognlto):

@amistre64

OpenStudy (amistre64):

2 sqr3 sinx = sec x 2 sqr3 sinx cosx = 1 2 sinx cosx = 1/sqrt3 sin(2x) = 1/sqrt3

OpenStudy (amistre64):

2x = arcsin(1/sqrt3) x = arcsin(1/sqrt3)/2 = .30773....

OpenStudy (lncognlto):

Thank you!

OpenStudy (amistre64):

does that make sense? or did i just do your work?

OpenStudy (lncognlto):

No, not at all. :) That does make sense. I thought there were supposed to be 2 answers to the question, so I was getting confused, and I kept the 2 apart from the sinxcosx. Now I remember 2 sinxcosx does equal sin2x.